OpenStudy (anonymous):

find the area of a triangle with sides of length 7 and 9 and included angle 72 degree.

OpenStudy (amistre64):

the height if going to have something to do with the sin of 72

OpenStudy (amistre64):

side1*side2*sin 72 / 2

OpenStudy (anonymous):



OpenStudy (amistre64):

is that really an included angle?

OpenStudy (anonymous):

the right angle no just the 72

OpenStudy (anonymous):

so 38.52

OpenStudy (amistre64):

an "included angle" is one the is nestled inbwteen the side lengths mentions

OpenStudy (amistre64):

 this is what you gave us in your information

OpenStudy (anonymous):

yea ... i drew it wrong i thought it was a triangle since it said find the area of a triangle

OpenStudy (amistre64):

:) then "height" can be regarded as 7sin(72) ; a base as 9 area tri = hb/2 = 9*7sin(72)/2

OpenStudy (anonymous):

so height 6.66

OpenStudy (anonymous):

wait know i am confused

OpenStudy (anonymous):

now*

OpenStudy (anonymous):

\[1/2 \times7\times9\times \sin 72 = 29.95828026 \approx 30\]

OpenStudy (anonymous):

so that is how you do it?

OpenStudy (anonymous):

yes .. half x 1st side x 2nd side x sine the angle included between them :)